### v2

@@ 1 Introduction
Early observations of the diffuse -ray emission at MeV scales [1, 2, 3] revealed the presence of a bright line at keV, from annihilations via para-positronium, that could not be explained solely by the positrons created by cosmic-ray (CR) collisions with the interstellar medium (ISM). In fact, these observations pointed to a continuous and steady positron injection from sources in the Galactic bulge. Recent measurements have detected the keV bulge emission with a significance exceeding and have identified a much more extended emission coinciding with the Galactic disk [4]. In addition, spectral analyses of the keV line have found high significance in the detection of continuum emission associated with ortho-positronium (the triplet state of positronium [5, 6]) and even in-flight positron annihilation (direct - annihilation of high energy positrons), revealing that positron sources contribute significantly to the diffuse -ray flux at MeV energies [7, 8].
Different sources have been proposed to explain the emission from the disk. Sources explaining contributions to the disk emission can be CR interactions, pulsars injecting , or sources synthesizing radioactive elements (like 26Al in massive stars, 24Ti in core-collapse supernovae or 56Ni in supernovae 1A) that decay into positrons ( decays) [9, 10]. Massive stars have received special attention since their emission can be observationally inferred and seem to be able to explain the full disk emission alone [11, 12, 13]. However, the measured bulge emission requires a spatial morphology and an injection rate that does not seem to easily fit with known candidates, such as low-mass X-ray binaries, type 1A supernovae or other sources expected to be located around the Galactic Center (GC) [12, 13, 10, 11, 14].
@@
Moreover, the calculations of the keV emission mentioned above did not take into account that positrons injected with energies above a few MeV can propagate appreciably from the injection source, thus smearing the expected keV profile with respect to the DM density profile [25]. In fact, Ref. [25] showed that a NFW profile is not able to explain the keV longitude profile when including propagation of the positrons with state-of-the-art propagation parameters. Ref. [25] also pointed out that the observed disk emission (and, in particular, at high longitudes) strongly constrains the DM production of , making it very difficult to explain the observed bulge emission (i.e. at the central longitudes) without exceeding the observations of the disk at high longitudes, and that a slope higher than would be necessary to explain the full keV line profile. This challenges the scenario where DM is the main culprit for bulge emission of the keV line: not only one needs to build models with an ad-hoc injection of neutrinos at early times, but one needs DM density profiles steeper than what suggested by numerical simulations.
@@ 2.1 DM distributions with a spike around Sgr A*
where is the Schwarzschild radius of the SMBH, and are the radial extension and the mass density profile of the DM spike respectively, and are the saturation radius and density of the spike due to DM annihilation respectively. is defined by , where
@@
The study of stellar orbits around Sgr A*, using the GRAVITY instrument since 2016 [45], has provided unprecedented precision in tracking these orbits. By analyzing astrometric and spectroscopic data from stars like S2, S29, S38, and S55, the collaboration has recently constrained any extended mass distribution around Sgr A* to be at most up to the S2 apocenter (which is around pc from the GC). This extended mass includes astrophysical objects (like white dwarfs and neutron stars) that, according to simulations, can contribute to it at a level close to the GRAVITY upper limit. The extended mass includes also a possible DM spike, thus potentially constraining our benchmarks. The cuspiest case considered in this work, i.e. the Gondolo-Silk spike model, has enclosed mass which easily complies with this constraint. The DM mass enclosed within pc from the GC, for a GS spike, would surpass a few hundreds for few pc, thus challenging the possibility of a much larger for the GS case (see footnote 2).
@@ 2.2 Positron injection and propagation
Given that our goal is to evaluate the diffuse keV emission as precisely as possible, we simulate the evolution of the position and energy of positrons with a state-of-the-art propagation set-up and approximate the thermalized distribution of positrons as the steady-state diffuse positron distribution generated from DM annihilations. In this way, we compute the keV line emission as proportional to the thermalized distribution of positrons. Following a similar procedure as in Ref. [48], we use a customized version [49] of the DRAGON2 code [50, 51], a dedicated CR propagation code prepared to simulate CR diffusion, accounting for all diffusion-reacceleration-advection-loss effects in the propagation of galactic CRs [52]. We use the same expression for the diffusion coefficient and propagation parameters (the best-fit values obtained from combined fits to AMS-02 CR data) as in Ref. [48], to which we refer the reader for more details. The injection per unit volume of particles from annihilating DM is given by the source term
@@
We remark that positrons with energies above the MeV-scale can propagate distances sufficienly far, before losing their energy and thermalizing, that their associated keV emission must be seen at a few degrees away from the GC. Especially, in the case where reacceleration is taken into account, MeV positrons could propagate hundreds of parsecs (see Fig. 9 in Appendix A). However, the conditions in the inner Galaxy are quite complex, in comparison to other Galactic regions, and the diffusion of charged particles may be significantly different. As an example, Ref. [54] found that MeV positrons could propagate kpc distances before thermalizing, because positrons would have negligible interactions with magnetohydrodynamic waves. We discuss the impact of the choice of different reacceleration benchmarks in App. A, and we anticipate that they do not affect the message of our paper.
@@ 2.3 Line emission
From the distribution of diffuse positrons in the Galaxy as a function of position, , we can calculate the emission of keV -rays, from the decay of para-positronium, in every point of the Galaxy. A complete evaluation should consider that the keV production is also proportional to the total electron density (either from free electrons and electrons bound to the different atomic and molecular gas species and to dust grains), and the cross section of positron annihilation (through positronium formation and direct annihilation) at the energy of the thermalized positrons (see, e.g. Ref. [55] for more details). Positrons are produced from DM annihilations with , but they thermalize with the medium before substantially annihilating with electrons and hence produce photons, see e.g. Refs. [7, 56]. With these considerations, the keV line flux integrated over the line of sight takes the following form
where is the fraction of positronium decays contributing to the keV line signal, is the energy-integrated flux of diffuse positrons () and the factor accounts for the emission of two keV photons per positron annihilation. is the solid angle element being , and the galactic longitude, latitude and distance from the Earth contributing to the signal. For the case of our interest of positrons sourced by DM annihilations, is obtained calculating the steady-state distribution of positrons obtained when propagating the injected positrons, given source term of Eq. (2.15), and including reacceleration and energy losses, as explained in Sec. 2.2. Analytical calculations of , directly from and the energy losses involved, can be obtained through Green functions [57, 58, 59, 60], although they are not valid when considering reacceleration, whose effect is very relevant at MeV energies [48].
@@
Given the uncertainties in the determination of the total electron density distribution (especially towards the GC) and since it is expected that it follows a smooth spatial distribution in the galactic disk without very high variations in the inner - kpc (roughly corresponding to the longitude range covered by the SPI data that we use here), we adopt the usual approach and estimate the distribution of the keV line emission as directly proportional to the spatial distribution of the diffuse positrons, thus setting the electron density to be uniform and with a value representative of the average electron density in the disk, cm-3 [61]. However, we expect that the electron density drops very quickly when moving out from the galactic plane, as the keV emission should as well. Off plane, the dominant gas phase is hot ionized gas. Therefore, we apply a scaling relation to the keV profiles, following the vertical distribution of free electron density in the Galaxy, also motivated by the fact that SPI observations of the latitude profile seem to show a similar variation of the keV latitude emission as that of the free electron density. We follow the same strategy as in Ref. [25] (where we refer the reader for more complete details), and include in our calculations the vertical dependence of the electron density from the NE2001 model [62, 63] of free electrons. Therefore, our predicted keV longitude profile follows the distribution of thermalized positrons in the galactic disk and the latitude profile follows the convolution of the vertical electron density distribution and the thermal positrons. We quantify in App. A the impact of using a different , we anticipate that doing so would further reinforce our conclusions.
@@
Then, as a conservative choice, we approximate the total annihilation cross sections as the well-known direct annihilation of positrons with electrons at rest [64] and assume that positrons annihilate once they reach the thermal energy of a warm medium (with K), as obtained in the analysis of the width of the keV line in Ref. [12]. This is a conservative estimation since we are not including charge-exchange production of positronium from when the positrons interact with electrons in atomic and molecular hydrogen (and helium), which may be more or less relevant depending on the medium. Taking a different thermal energy would lead to change in the normalization of the signal, meaning our conclusions (except those on the values of best-fit DM annihilation cross sections) would remain unchanged.
@@ 2.4 Continuum emission
Ortho-positronium (o-ps) emission. We employ the spectral function derived by Ore and Powell 1949 [67], which is also used by the SPI collaboration. The o-positronium emission spectrum then reads
@@
with keV being the mass of the electron in natural units. It is normalized (i.e. we set ) such that the integrated emission in a given region satisfies , with , as was found from the analysis of Ref. [68]. Theoretically, the value of this ratio can be at most, but can also take lower values. Among all the continuum emissions from DM, the one from orthopositronium dominates at photon energies below 511 keV.
@@
In-flight annihilation emission. We use the same prescription as in Beacom and Yuksel [7] with a modification to account for the fact that injected positrons are not monoenergetic, instead having a distribution of energies. First, we calculate the average keV flux per solid angle, as detailed above, in the region of the SPI continuum data that we use (, ) and then we use the following expression to evaluate the -ray flux from in-flight annihilations [56]
@@
where the second integral over accounts for the energy-distribution of injected positrons determined by , being bounded between the highest energy used in the positron propagation simulation ( GeV), and the minimum positron energy able to produce -rays with energy . The term represents the fraction of positrons at the energy contributing to the 511 keV emission (in the limit where this fraction is unity, we recover the case of mono-energetic positrons generating the total 511 keV flux). This takes into account the energy loss for a positron of energy to the final energy before annihilating; the number density of hydrogen, , counts the number of targets on which positrons scatter and lose energy. is the probability, per unit of energy (taken from Eq. (4) of Ref. [7]), for a positron with initial energy to produce a ray in flight before reaching the energy ; similarly, is the probability for a positron with initial energy to produce a photon before thermalizing, and [68] is the fraction of positrons annihilating through positronium states. We notice that this expression eventually becomes independent of the gas density, , since (taken to be the ionization energy losses from the DRAGON2 code, including H and He – see Ref. [50] for more details) is directly proportional to .
@@
The FSR emission is lower than the in-flight annihilation emission by a factor of a few and also lower than the o-ps emission at sub-MeV energies, as expected from previous calculations (see, e.g. Ref. [8]).
@@
Background emission. On top of the DM-induced -ray continuum signals, one has to account for two important backgrounds: the extragalactic background light and the galactic IC emission which is the dominant source here. For the galactic component, we use the electron model from Refs. [72, 73], which is optimized to reproduce the electron and positron emission at the Earth location, as well as the local -ray emissivity down to a few tens of MeV. We have checked that this model shows very good agreement with the data down to keV. Since the SPI instrument is not sensitive to isotropic backgrounds (since they are degenerated with instrumental noise) we do not add any isotropic extragalactic component. Furthermore, the extragalactic background emission is very low in comparison to the other positron-induced emission components.
@@ 3 Comparison with INTEGRAL-SPI data
To illustrate the importance of accounting for a spike in the DM distribution, we show in Fig. 3 a comparison of the expected longitudinal profile of the keV emission for a DM mass of MeV, for four different DM density profiles reviewed in Sec. 2.1: NFW [17], *MIN and *MAX [27] and Gondolo-Silk [76]. In order to be consistent with the measured disk emission (which is likely dominated by emission from astrophysical sources), these are normalized to an annihilation cross section, , that ensures we do not overshoot any high-longitude data point (except for the one at , which is, in every case, exceeded by less than ).
@@
Still, in both cases there is an apparent residual asymmetry in the data, showing a slightly higher flux at keV towards right longitudes that we cannot reproduce. In fact, the SPI collaboration found that “when the disk is separated into an eastern and a western hemisphere, the line widths from positive and negative longitudes also show a discrepancy at the level” [4]. At the moment, it seems difficult to ascertain whether this asymmetry is just an artifact coming from systematic sources of uncertainty or it has a physical origin. The current analyses from the SPI collaboration seem to reproduce the asymmetric data [4], when they tested templates with a line that is slightly displaced from the GC (see discussions in Ref. [77, 74, 78, 79, 80]). In any case, we note that this asymmetry cannot be explained by DM and we would need an extra asymmetric component to account for it. A population of sources located such that they follow the spiral arms could produce an asymmetric disk emission (see Fig. 6 of Ref. [81]), however, given the low significance of the asymmetry, we do not perform any detailed analysis of this for now. We note that the stellar bulge template use in Ref. [9] would predict such a feature.
For completeness, we also show in Fig. 5 the contribution of the disk and DM signals separately, for the latitudinal keV emission profile. As we see, the contribution of the disk to the central latitudes can be as important as a few tens of percent of the total measured emission. This leads to the fact that including the disk contribution should reduce the DM-induced signals (both, line-like and continuum) and relax previous constraints on DM models trying to explain the bulge emission. Concretely, with our best-fit models shown in Fig. 4 we obtain that the disk emission have a contribution to the total keV emission in the inner ( in the inner ) for the best-fit Gondolo-Silk signal at MeV mass, while for a MeV mass this contribution is of in the inner ( in the inner ). Similarly, for the *MIN profile, the disk contribution is of in the inner ( in the inner ) for a MeV DM mass and of in the inner ( in the inner ) for a MeV mass.

### v3

@@ 1 Introduction
Early observations of the diffuse -ray emission at MeV scales [1, 2, 3] revealed the presence of a bright line at keV, from annihilations via para-positronium, that could not be explained solely by the positrons created by cosmic-ray (CR) collisions with the interstellar medium (ISM). In fact, these observations pointed to a continuous positron injection from sources in the Galactic bulge. Recent measurements have detected the keV bulge emission with a significance exceeding and have identified a much more extended emission coinciding with the Galactic disk [4]. In addition, spectral analyses of the keV line have found high significance in the detection of continuum emission associated with ortho-positronium (the triplet state of positronium [5, 6]) and even in-flight positron annihilation (direct - annihilation of high energy positrons), revealing that positron sources contribute significantly to the diffuse -ray flux at MeV energies [7, 8].
Different sources have been proposed to explain the emission from the disk. Sources explaining contributions to the disk emission can be CR interactions, pulsars injecting , or sources synthesizing radioactive elements (like 26Al in massive stars, 44Ti in core-collapse supernovae or 56Ni in supernovae 1A) that decay into positrons ( decays) [9, 10]. Massive stars have received special attention since their emission can be observationally inferred and could be able to explain a significant portion of the disk emission [11, 12, 13]. However, the measured bulge emission requires a spatial morphology and an injection rate that does not seem to easily fit with known candidates, such as low-mass X-ray binaries, type 1A supernovae or other sources expected to be located around the Galactic Center (GC) [12, 13, 10, 11, 14].
@@
Moreover, the calculations of the keV emission mentioned above did not take into account that positrons injected with energies above a few MeV can propagate appreciably from the injection source, thus smearing the expected keV profile with respect to the DM density profile [25]. In fact, Ref. [25] showed that a NFW profile is not able to explain the keV longitude profile when including propagation of the positrons with state-of-the-art propagation parameters. Ref. [25] also pointed out that the observed disk emission (and, in particular, at high longitudes) strongly constrains the DM production of , making it very difficult to explain the observed bulge emission (i.e. at the central longitudes) without exceeding the observations of the disk at high longitudes, and that when including the effects of propagation, in a similar way as we do in this work, a slope around would be necessary to explain the full keV line profile (however, if no broadening of the signal by diffusion is considered, a profile with notably reproduces the morphology of the line). This challenges the scenario where DM is the main culprit for bulge emission of the keV line: not only one needs to build models with an ad-hoc injection of neutrinos at early times, but one needs DM density profiles steeper than what suggested by numerical simulations.
@@ 2.1 DM distributions with a spike around Sgr A*
where is the Schwarzschild radius of the supermassive black hole (SMBH), and are the radial extension and the mass density profile of the DM spike respectively, and are the saturation radius and density of the spike due to DM annihilation respectively. is defined by , where
@@
The study of stellar orbits around Sgr A*, using the GRAVITY instrument since 2016 [45], has provided unprecedented precision in tracking these orbits. By analyzing astrometric and spectroscopic data from stars like S2, S29, S38, and S55, the collaboration has recently constrained any extended mass distribution around Sgr A* to be at most up to the S2 apocenter (which is around pc from the GC). This extended mass includes astrophysical objects (like white dwarfs and neutron stars) that, according to simulations, can contribute to it at a level close to the GRAVITY upper limit. The extended mass includes also a possible DM spike, thus potentially constraining our benchmarks. The cuspiest case considered in this work, i.e. the Gondolo-Silk spike model, has enclosed mass which easily complies with this constraint. The DM mass enclosed within pc from the GC, for a GS spike, would surpass a few hundreds for few pc (in particular, for our choice of DM density profiles, the enclosed mass within this region is 0.129287, 26.941, 1.9746, 1.00483 M⊙ for the an NFW, Gondolo-Silk, *Min and *Max, respectively), thus challenging the possibility of a much larger for the GS case (see footnote 2).
@@ 2.2 Positron injection and propagation
Given that our goal is to evaluate the diffuse keV emission as precisely as possible, we simulate the evolution of the position and energy of positrons with a state-of-the-art propagation set-up and approximate the thermalized distribution of positrons as the steady-state diffuse positron distribution generated from DM annihilations. In this way, we compute the keV line emission as proportional to the thermalized distribution of positrons. Following a similar procedure as in Ref. [48], we use a customized version [49] of the DRAGON2 code [50, 51], a dedicated CR propagation code prepared to simulate CR diffusion, accounting for all diffusion-reacceleration-advection-loss effects in the propagation of galactic CRs [52]. We use the same expression for the diffusion coefficient and propagation parameters (the best-fit values obtained from combined fits to AMS-02 CR data) as in Ref. [48], to which we refer the reader for more details. The injection per unit volume of particles from annihilating self-conjugate DM is given by the source term
@@
The propagation of positrons at sub-MeV energies remains poorly understood, particularly in the sub-keV range. At these low energies, the standard diffusion model may break down, potentially preventing positrons from propagating over significant distances. Additionally, if positrons experience nearly catastrophic energy losses or if diffusion mechanisms operate in an unconventional manner, their propagation could be severely limited, possibly confining them to scales below a parsec. These uncertainties make it challenging to reliably model positron behavior at sub-MeV energies, which is why such cases are often excluded from consideration. However, for the sake of having a reasonable estimate, we use an extrapolation of the diffusion coefficient that is set from CR observations at higher energies. Under this assumption, positrons with energies above the MeV-scale can propagate distances sufficiently far, before losing their energy and thermalizing, that their associated keV emission must be seen at a few degrees away from the GC. Especially, in the case where reacceleration is taken into account, MeV positrons could propagate hundreds of parsecs (see Fig. 9 in Appendix A). However, the conditions in the inner Galaxy are quite complex, in comparison to other Galactic regions, and the diffusion of charged particles may be significantly different. As an example, Ref. [54] found that MeV positrons could propagate kpc distances before thermalizing, because positrons would have negligible interactions with magnetohydrodynamic waves. We discuss the impact of the choice of different reacceleration benchmarks in App. A, and we anticipate that they do not affect the message of our paper.
@@ 2.3 Line emission
From the distribution of diffuse positrons in the Galaxy as a function of position, , we can calculate the emission of keV -rays, from the decay of para-positronium, in every point of the Galaxy. A complete evaluation should consider that the keV production is also proportional to the total electron density (either from free electrons and electrons bound to the different atomic and molecular gas species and to dust grains), and the cross section of positron annihilation (through positronium formation from charge-exchange with hydrogen gas, which is expected to be the dominant mechanism of positronium production) at the energy of the thermalized positrons (see, e.g. Ref. [55] for more details). Positrons are produced from DM annihilations with , but they thermalize with the medium before substantially annihilating with electrons and hence produce photons, see e.g. Refs. [7, 56]. With these considerations, the keV line flux integrated over the line of sight takes the following form
where is the fraction of positronium decays contributing to the keV line signal, is the steady-state flux of diffuse positrons at the thermal energy and the factor accounts for the emission of two keV photons per positron annihilation. is the solid angle element being , and the galactic longitude, latitude and distance from the Earth contributing to the signal. We note that an additional component the direct annihilation of positrons with electrons, which also contributes to the total line emission. However, the fraction of positrons annihilating before reaching positronium state is found to be very small ( [57] - see more details in Sect. 2.4 too).
@@
Given the uncertainties in the determination of the total electron density distribution (especially towards the GC) and since it is expected that it follows a smooth spatial distribution in the galactic disk without very high variations in the inner - kpc (roughly corresponding to the longitude range covered by the SPI data that we use here), we adopt the usual approach and estimate the distribution of the keV line emission as directly proportional to the spatial distribution of the diffuse positrons, thus setting the electron density to be uniform and with a value representative of the average electron density in the disk, cm-3 [58]. However, we expect that the electron density drops very quickly when moving out from the galactic plane, as the keV emission should as well. Observations by SPI have shown that the morphology of the 511 keV line favours the formation of positronium in an even mixture of warm hydrogen gas and electron gas. To capture this, we apply a scaling relation to the keV profiles, following the vertical distribution of an equal sum of warm gas and electron density in the Galaxy. In particular, for our benchmark evaluation, we implement a 50% Nakanishi [59] (representing moderately warm gas medium) plus a 50% Ferriere [60] (for the electron density) to be consistent with these observations. Meanwhile, we adopt a constant electron density in the Galactic plane, as discussed above. In this way, our predicted keV longitude profile follows the distribution of thermalised positrons in the Galactic disk and the latitude profile follows the convolution of the vertical electron density distribution and the thermal positrons. We show the effects of convolving with the different radial gas distributions in the right panel of Fig. 5, and discuss it in more detail in App. A.
@@
Then, we approximate the total annihilation cross sections, , as the charge-exchange cross section with hydrogen [61], that is about a million times greater than the direct annihilation of positrons with free electrons [62], and assume that positrons annihilate once they reach the thermal energy of a warm medium (with K), as obtained in the analysis of the width of the keV line in Ref. [57]. Note that this implies that of Eq. (2.16) is dominated by the electrons in hydrogen for the purpose of computing the 511 keV signal. Taking a different thermal energy would lead to change in the normalization of the signal, meaning our conclusions (except those on the values of best-fit DM annihilation cross sections) would remain unchanged.
@@ 2.4 Continuum emission
Ortho-positronium (o-ps) emission. We employ the spectral function derived by Ore and Powell 1949 [65], which is also used by the SPI publications. The o-positronium emission spectrum then reads
@@
with keV being the mass of the electron in natural units. It is normalized (i.e. we set ) such that the integrated emission in a given region satisfies , with , as was found from the analysis of Ref. [57]. Theoretically, the value of this ratio can be at most, but can also take lower values down to 0. Among all the continuum emissions from DM, the one from orthopositronium dominates at photon energies below 511 keV.
@@
In-flight annihilation emission. We use the same prescription as in Beacom and Yuksel [7] with a modification to account for the fact that injected positrons are not monoenergetic, instead having a distribution of energies (as in Ref. [66]). First, we calculate the average keV flux per solid angle, as detailed above, in the region of the SPI continuum data that we use (, ) and then we use the following expression to evaluate the -ray flux from in-flight annihilations [56]. 444This implies a simplifying assumption that in-flight annihilation occurs, on average, within the same region as the thermalised positrons. It is important to note that in-flight annihilations may, in reality, occur in different regions depending on the propagation of positrons. A more accurate estimation would require a detailed propagation model, but for the sake of simplicity, we proceed with this assumption.
@@
where the second integral over accounts for the energy-distribution of injected positrons determined by , being bounded between the highest energy used in the positron propagation simulation ( GeV), and the minimum positron energy able to produce -rays with energy . The lower bound for the photon energy is half the electron mass. The term represents the fraction of positrons at the energy contributing to the 511 keV emission (in the limit where this fraction is unity, we recover the case of mono-energetic positrons generating the total 511 keV flux). This takes into account the energy loss for a positron of energy to the final energy before annihilating; the number density of hydrogen, , counts the number of targets on which positrons scatter and lose energy. is the probability, per unit of energy (taken from Eq. (4) of Ref. [7]), for a positron with initial energy to produce a ray in flight before reaching the energy ; similarly, is the probability for a positron with initial energy to produce a photon before thermalizing, and [57] is the fraction of positrons annihilating through positronium states. We notice that this expression eventually becomes independent of the gas density, , since (taken to be the ionization energy losses from the DRAGON2 code, including H and He – see Ref. [50] for more details) is directly proportional to . However, it is important to emphasize that this proportionality does not hold universally. For instance, the cross sections for interactions with H and He have different dependencies and weightings compared to energy losses, and the target density for direct annihilation is typically the electron density rather than the hydrogen density. Additionally, at energies above the positronium formation threshold, the direct annihilation cross section with electrons from H or He is not well characterized, though it is known to be 1-2 orders of magnitude smaller than with free electrons at lower energies. For GeV positrons, energy losses such as Inverse Compton scattering and synchrotron radiation, which do not depend on the H and He number density, become significant and do not cancel out. Thus, the assumption of independence from is valid within the specific context of sub-MeV propagation in a single-atom medium which we consider here.
@@
The FSR emission is lower than the in-flight annihilation emission by a factor of a few and also lower than the o-ps emission over the relevant energies ( keV), as expected from previous calculations (see, e.g. Ref. [8]).
@@
Background emission. On top of the DM-induced -ray continuum signals, one has to account for two important backgrounds: the extragalactic background light and the galactic IC emission which is the dominant source here. For the galactic component, we use the electron model from Refs. [70, 71], which is optimized to reproduce the electron and positron emission at the Earth location, as well as the local -ray emissivity down to a few tens of MeV. We have checked that this model shows very good agreement with the data down to keV. The SPI instrument, which uses a coded mask for contrast measurements, is not sensitive to isotropic backgrounds. This is because isotropic backgrounds do not produce a detectable shadow pattern and are therefore degenerate with the instrumental noise.
@@ 3 Comparison with INTEGRAL-SPI data
To illustrate the importance of accounting for a spike in the DM distribution, we show in Fig. 3 a comparison of the expected longitudinal profile of the keV emission for a DM mass of MeV, for four different DM density profiles reviewed in Sec. 2.1: NFW [17], *MIN and *MAX [27] and Gondolo-Silk [74]. In this comparison, our lines are the result of integrating the emission per unit of solid angle over , which is approximately SPI’s spatial resolution. In order to be consistent with the measured disk emission (which is likely dominated by emission from astrophysical sources), these are normalized to an annihilation cross section, , that ensures we do not overshoot any high-longitude data point (except for the one at , which is, in every case, exceeded by less than ).
@@
Moreover, Ref. [4] found that “when the disk is separated into an eastern and a western hemisphere, the line widths from positive and negative longitudes also show a discrepancy at the level”. At the moment, it seems difficult to ascertain whether this asymmetry is just an artifact coming from systematic sources of uncertainty or it has a physical origin. However, asymmetries in the fluxes seem to be not significant, as Ref. [72] reported. Current analyses seem to prefer symmetric fluxes [4] when templates with a line that is slightly displaced from the GC is added (see discussions in Ref. [75, 72, 76, 77, 78]). In any case, it would be challenging to explain, via DM, this asymmetry or displacement of the peak emission with respect to the GC. Even if some hydrodynamic simulations of Milky-Way-like galaxies find offsets between the central DM peak and the GC (see e.g. Ref. [79]), they come with a flat DM central distribution and/or could affect the DM spike. Further study is then needed to determine whether or not they could explain the asymmetric emission in our context. A population of sources located such that they follow the spiral arms could produce an asymmetric disk emission (see Fig. 6 of Ref. [80]), however, given the low significance of the asymmetry, we do not perform any detailed analysis of this for now. We note that the stellar bulge template use in Ref. [9] would predict such a feature too.
For completeness, we also show in the left panel of Fig. 5 the contribution of the disk and DM signals separately, for the latitudinal keV emission profile. As we see, the contribution of the disk to the central latitudes can be as important as a few tens of percent of the total measured emission. This leads to the fact that including the disk contribution should reduce the DM-induced signals (both, line-like and continuum) and relax previous constraints on DM models trying to explain the bulge emission. Concretely, with our best-fit models shown in Fig. 4 we obtain that the disk emission have a contribution to the total keV emission in the inner ( in the inner ) for the best-fit Gondolo-Silk signal at MeV mass, while for a MeV mass this contribution is of in the inner ( in the inner ). Similarly, for the *MIN profile, the disk contribution is of in the inner ( in the inner ) for a MeV DM mass and of in the inner ( in the inner ) for a MeV mass. Additionally, in the right panel of Fig. 5, we show the longitudinal profile of the 511 keV emission expected from DM (assuming a Gondolo-Silk profile) convolved with different radial gas distributions: The Nakanishi [59], Ferriere [60] and an even combination of Nakanishi and Ferrire distributions (“Combined dist” in the legend). As explained in Sect 2.3, our benchmark case follows a constant gas density in the Galactic plane, convolved with the vertical distribution of electrons expected from the combination of the Nakanishi and Ferriere models. Therefore, the longitudinal profile in our benchmark scenario directly features the distribution of thermal positrons in the disk. See more details in Appendix A.